OpenStudy (anonymous):

In a bowl of jelly beans there are 12 red, 10 black, 5 white, and 3 green. If three jellly beans are chosen one at a time, what is the probability that 3 black jelly beans will be chosen? WILL MEDAL

OpenStudy (anonymous):

\[ \frac{10}{12+10+5+3}\cdot \frac{9}{12+9+5+3}\cdot \frac{8}{12+8+5+3} \]

OpenStudy (kropot72):

This is sampling without replacement: \[\large P(3\ black)=\frac{10C3}{30C3}=\frac{10\times9\times8}{30\times29\times28}=you\ can\ calculate\]

OpenStudy (anonymous):

\[ ^{10}P_3 = \frac{10!}{(10-3)!} = 10\times 9\times 8 \]